• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 00:21:56 on June 24, 2010

    Entry number 328103

    keyword=[APEX] Experiment status

    We did not have time to have the lead glass put in today, this will be
    done tomorrow morning. Owl shifts are canceled.
    
    The VDC amplifier/dsicriminator card thresholds were checked, the results
    of the checks will be recorded in another HALOG entry by the people who
    did the test.
    
    Sergey made a S0 trigger fed into the T1 input of the trigger supervisor.
    
    The RG213 cables for the gas cerenkov were moved from the testlab and
    checked and are ready to be used.
    
    Sergey is taking data and will start checking the VDC with cosmics as
    soon as he generates the updated crate map for the VDC.
    
    Main things left to be done for tomorrow :
    - finish single arm / coincidence trigger
    - install / cable pion rejector
    - cable gas cerenkov
    - if radiation level are low enough change the O-ring